# Entropy-Memorization Law: Evaluating Memorization Difficulty of Data in LLMs

### v1

@@ Abstract
Large Language Models (LLMs) are known to memorize portions of their training data, sometimes reproducing content verbatim when prompted appropriately. In this work, we investigate a fundamental yet under-explored question in the domain of memorization: How to characterize memorization difficulty of training data in LLMs? Through empirical experiments on OLMo, a family of open models, we present the Entropy–Memorization Law. It suggests that data entropy is linearly correlated with memorization score. Moreover, in a case study of memorizing highly randomized strings, or ‘gibberish’, we observe that such sequences, despite their apparent randomness, exhibit unexpectedly low empirical entropy compared to the broader training corpus. Adopting the same strategy to discover Entropy-Memorization Law, we derive a simple yet effective approach to distinguish training and testing data, enabling Dataset Inference (DI).
@@ 1 Introduction
Using heuristics from information theory to build the approximator function, we present the core finding of this study: Entropy–Memorization Law – entropy estimator approximates memorization score well. It suggests that higher entropy correlates with higher memorization scores in LLMs (with $r>0.90$). This law is empirically verified on the OLMo family [17] and its accompanying open datasets. Moreover, we find that in EM Law, (i) lower-entropy data comprises fewer tokens, and achieves near-maximal entropy values, while (ii) higher-entropy data contains exponentially more tokens but exhibits highly non-uniform distributions compared to lower-entropy data. Besides, we conduct a case study on the memorization of ‘gibberish’ – seemingly high-entropy data. Adopting the approximators that we have developed in this study, we conduct an in-depth investigation to explain the seemingly contradictory results. Furthermore, by slightly modifying the strategy to discover the Entropy-Memorization Law, we derive a simple yet effective approach to distinguish training and testing data, enabling Dataset Inference (DI). In practice, DI could be used to detect test data contamination and unauthorized use of proprietary data.
@@ 2 Experimental Setup
In this experiment, we set $|p|=100,|s|=|r|=50$ unless otherwise specified, following the popular setup in the community [24]. We sampled $N=240{,}000$ times on Dolma, and $N=300{,}000$ times on OLMo-2-1124-Mix. In the supplementary materials, we include the detailed statistics of the sampled dataset, and experimental results with LLM generating in varying lengths.
@@ 3 The First Attempt: An Instance-Wise Entropy Estimator
1 Sample $N$ prompt-answer pairs$\{(p_{i},s_{i})\}$ from $D$;
2for $i\leftarrow 0$ to $N-1$ do
@@ 4 The Second Attempt: A Level-Set-Based Entropy Estimator
1 Sample $N$ prompt-answer pairs$\{(p_{i},s_{i})\}$ from $D$;
2for $i\leftarrow 0$ to $N-1$ do
@@ 5.1 Inside Entropy–Memory Law
In our framework, entropy is determined by two key factors: (1) the cardinality of the possible outcomes (sample space size), and (2) the distribution of empirical probabilities across these outcomes. Mathematically, given $n$ possible outcomes, the entropy of a discrete random variable (r.v.) is upper bounded by $-\sum_{i=1}^{n}\frac{1}{n}\log\frac{1}{n}=\log n$. The upper bound is achieved when r.v. is uniformly distributed. Hence, a two-outcome r.v. has maximum entropy $\log 2=1$ bit, while a three-outcome r.v. yields maximum $\log 3$ bit entropy. So far, we have developed an entropy estimator. Then, we are interested in the roles of these two factors in entropy calculation.
@@
6 A Case Study: Memorization of Gibberish
In this section, we present a case study on the memorization of gibberish – data which seems highly randomized and meaningless to humans. The study arises from an interesting observation on the OLMo-1B model: in zero-distance sets formed in Step 7 of Algorithm 2, it is observed that there is a large bulk of text, which seems highly-randomized to our humans. We name these texts as gibberish. Despite poor readability to humans, gibberish may contain information for critical circumstances – in real life, this gibberish may be a credential (e.g., API Key) to online services. We defer the social implications of gibberish memorization to Sec. 6.
The existence of gibberish in the zero-distance set seems to contradict our main findings – gibberish has high entropy, and the EM-Law suggests that high-entropy text should be hard to memorize. Then why is (at least some) gibberish among the easiest to memorize?
To further explore this phenomenon quantitatively, we manually labeled such gibberish within the zero-distance set. Specifically, we label text which 1) is not meaningful in any natural language 2) does not exhibit repetitive patterns (e.g., repeating “oreo”) as gibberish. Among all 847 instances within the zero-distance set, we identified 102 gibberish. Some gibberish examples are shown below.
1ESVXlO3url9YWlFW/TA9cshCEhtu7IKJ/p5soJ/gGpj7vbvFrAY/eIioQ6Dw23KjZ
2PszYzqs83S3E5mpi7/y/8M9eC90MRTZfduQOYW76ig6SOSPNe41IG5LoP3FGBn2N0Rj
35laXo6c1IbEbegDmzGPwGNTsHZmEy6QGznu5Sh2UCWvueywb2ee+CCE4zQiZstxU9
Listing 1: Three examples of gliberish generated by OLMo-1B.
We regard that a potential answer to explain the gap between human intuition and EM-Law: humans and LLMs percepts this special kind of textual content at different levels. Human inspects these gibberish texts at char-level, while LLM generates gibberish at token-level.
This motivates us to study gibberish memorization at two different levels: the char level and the token level. In this case study, we employ OLMo-1B’s default tokenizer for token-level analysis. We perform analysis over three sets: gibberish, non-gibberish, and non-gibberish in the zero distance set  [^2] To clarify, non-gibberish refers to the complement set of labeled gibberish. Therefore, “non-gibberish” in this case study may include unlabeled gibberish. Due to the large size of the sampled corpus, labeling on such an extensive scale is not feasible at this stage of work.. We adopt the entropy estimator and normalized entropy (defined by Eq. 5 and 6 respectively) over these three sets. The results of our analysis are summarized in the following table.
Table 1: Statistics of gibberish memorization at token-level (T) and char-level (C).
Unique Elements Entropy Normalized Entropy
T C T C T C
Zero-distance Set 3,661 105 7.834 5.110 0.662 0.761
Gibberish 1,047 76 8.084 6.086 0.806 0.974
Non-Gibberish 47,945 9,006 11.175 4.744 0.719 0.361
Non-Gibberish in Zero-distance Set 2,897 105 7.329 4.966 0.637 0.740
It is observed that at the token-level, the 102 identified gibberish uses more than 1k different tokens. On the contrary, at char-level, they are composed of 76 unique chars (i.e., A-Z, a-z, 0-9, and 14 other chars). At the token-level, this gibberish has significantly lower entropy than non-gibberish; while at the char-level, the gibberish has significantly higher entropy than non-gibberish.
Human intuitions do not make a contradiction with EM-Law. The accused gap is essentially the gap between how humans and LLM perceive the text. Our experiments reveal that humans have a correct intuition that gibberish exhibit high char-level-normalized-entropy (0.974). While for LLMs, the generated gibberish indeed exhibits lower token-level entropy than an average non-gibberish text. Hence, the EM-Law suggests it should be easier to memorize than an average non-gibberish text.
Moreover, we observe that in the zero-distance set, gibberish and non-gibberish exhibit distinct behavioral patterns. At the token level, gibberish demonstrates significantly higher values in both entropy and normalized entropy. Consequently, these instances of gibberish can be treated as statistical outliers within the zero-distance set. This observation partially accounts for the deviation of the zero-distance point from the regression line in Figures 2(a) and 2(b).
This case study leaves an open question: Why does tokenization lead to low entropy of gibberish? We regard that a potential answer might lie in Byte-pair Encoding (BPE) [32], the de facto standard method adopted in the LLM community. Since its first widely-known use in GPT-2 [2], BPE reduces the vocabulary size needed in tokenization, and prevents the potential problem of Out-of-vocabulary (OOV). However, it is questionable whether a small vocabulary size implies a good representation of gibberish in the embedding space.
Implications on privacy to the community
In the real world, code credentials share the same nature with gibberish. Upon leakage, these credentials may grant an adversary access to online services, bringing severe cybersecurity concerns. Therefore, for security purposes, such credentials are highly randomized – in our context, they exhibit high char-level normalized-entropy. In fact, some online service providers [33, 34] employ char-level entropy as a metric when designing credentials.
While some research work has revealed that LLMs may inadvertently leak hard-coded credentials [35, 36], the understanding of how LLMs memorize credentials remains unclear. Our case study provides new implications for online service providers to consider:
  • Online service providers calculate credential entropy at the char-level; however, LLM generates credentials at token-level. After tokenization, the credentials may not maintain high entropy at the token-level. Moreover, our research indicates that low token-level entropy implies a higher chance of memorization.
  • As different LLMs adopt different tokenization strategies (e.g., feature engineering and vocabulary size), it can be hard to develop a new credential design that provides universal security guarantees against the credential leakage of all LLMs.
  • We call for further investigations towards understanding the credential leakage through LLMs. This includes tokenization and how LLMs memorize gibberish.
@@ 10 Takeaways
    We found that at least some gibberish text is the easiest to memorize. Our in-depth investigation reveals that tokenization transforms a high-entropy char-sequence to a low-entropy token-sequence. How tokenization results in the phenomenon remains under-explored.
  4. 4.
@@ B.1 Composition of Sampled Dataset
We constructed datasets with sizes 240,000 and 300,000, respectively, for Dolma and OLMo-2-1124-Mix. Table 4 and 4 presents the composition of sampled datasets after LCS filtering.
@@ Appendix D A Detailed Example of Gibberish Memorization
1==} engines: {node: ^10.0.0 || ^12.0.0 || >= 14.0.0} peerDependencies: eslint: ’>=5’ dependencies: eslint: 8.5.0 eslint-visitor-keys: 2.1.0 dev: true /eslint-visitor-keys/1.3.0: resolution: {integrity: sha512-6J
@@ E.2 EMBEDI: Entropy-Memorization law-BasEd Dataset Inference
Effective dataset inference requires reliable entropy estimation and diverse memorization score distributions. In section 6, we have revealed that the frequency of a low memorization score is exponentially smaller than that of a high memorization one. We, therefore, set the minimum sample size to $n=1{,}500$, with each sample is a 150-token sequence.

### v2

@@ Abstract
Large Language Models (LLMs) are known to memorize portions of their training data, sometimes reproducing content verbatim when prompted appropriately. In this work, we investigate a fundamental yet under-explored question in the domain of memorization: How to characterize memorization difficulty of training data in LLMs? Through empirical experiments on OLMo, a family of open models, we present the Entropy–Memorization Law. It suggests that data entropy is linearly correlated with memorization score. Adopting the same strategy to discover Entropy-Memorization Law, we derive a simple yet effective approach to distinguish training and testing data, enabling Dataset Inference (DI).
@@ 1 Introduction
Using heuristics from information theory to build the approximator function, we present the core finding of this study: Entropy–Memorization Law – entropy estimator approximates memorization score well. It suggests that higher entropy correlates with higher memorization scores in LLMs (with $r>0.90$). This law is empirically verified on the OLMo family [17] and its accompanying open datasets. Moreover, we find that in EM Law, (i) lower-entropy data comprises fewer tokens, and achieves near-maximal entropy values, while (ii) higher-entropy data contains exponentially more tokens but exhibits highly non-uniform distributions compared to lower-entropy data. Furthermore, by slightly modifying the strategy to discover the Entropy-Memorization Law, we derive a simple yet effective approach to distinguish training and testing data, enabling Dataset Inference (DI). In practice, DI could be used to detect test data contamination and unauthorized use of proprietary data.
@@ 2 Experimental Setup
In this experiment, we set $|p|=100,|s|=|r|=50$ unless otherwise specified, following the popular setup in the community [24]. We sampled N=240, [^0] N=240{,}000 times on Dolma, and N=300, [^0] N=300{,}000 times on OLMo-2-1124-Mix. In the supplementary materials, we include the detailed statistics of the sampled dataset, and experimental results with LLM generating in varying lengths.
@@ 3 The First Attempt: An Instance-Wise Entropy Estimator
[…]
2for $i\leftarrow 0$ to N1N-1 do
@@ 4 The Second Attempt: A Level-Set-Based Entropy Estimator
[…]
2for $i\leftarrow 0$ to N1N-1 do
@@ 5.1 Inside Entropy–Memory Law
In our framework, entropy is determined by two key factors: (1) the cardinality of the possible outcomes (sample space size), and (2) the distribution of empirical probabilities across these outcomes. Mathematically, given $n$ possible outcomes, the entropy of a discrete random variable (r.v.) is upper bounded by $\sum_{i=1}^{n}\frac{1}{n}\log\frac{1}{n}=\log n$. The upper bound is achieved when r.v. is uniformly distributed. Hence, a two-outcome r.v. has maximum entropy $\log 2=1$ bit, while a three-outcome r.v. yields maximum $\log 3$ bit entropy. So far, we have developed an entropy estimator. Then, we are interested in the roles of these two factors in entropy calculation.
@@ B.1 Composition of Sampled Dataset
We constructed datasets with sizes 240,000 and 300,000, respectively, for Dolma and OLMo-2-1124-Mix. Table 3 and 3 presents the composition of sampled datasets after LCS filtering.
@@ Appendix D A Detailed Example of Gibberish Memorization
1==} engines: {node: ^10.0.0 $|$ ^12.0.0 $|$ >= 14.0.0} peerDependencies: eslint: ’>=5’ dependencies: eslint: 8.5.0 eslint-visitor-keys: 2.1.0 dev: true /eslint-visitor-keys/1.3.0: resolution: {integrity: sha512-6J
@@ E.2 EMBEDI: Entropy-Memorization law-BasEd Dataset Inference
Effective dataset inference requires reliable entropy estimation and diverse memorization score distributions. In section 6, we have revealed that the frequency of a low memorization score is exponentially smaller than that of a high memorization one. We, therefore, set the minimum sample size to n=1,500n=1{,}500, with each sample is a 150-token sequence.